# Depth-13 Sorting Networks for 28 Channels

### v1

@@ 1 Introduction
Our construction of n=28 and d=13 is based on ehlers2017merging. As we have more channels and one more layer, we use two new techniques to reduce the search space: assume the network is reflection-symmetric, and greedily extend the 24-channel prefix by one layer before solving the SAT problem.
@@ 2 Preliminaries and Redundant Prefixes
We formalize some notation. For an integer $n$, we use $[n]$ to denote the set $\{0,1,\ldots,n-1\}$. A comparator is an unordered pair of distinct channels, denoted as $(i,j)$ where $i<j$. A layer is a set of comparators that can be applied in parallel, meaning no two comparators share a common channel.
@@
If a comparison network with $n$ inputs sorts all $2^{n}$ Boolean sequences into non-decreasing order, then it correctly sorts all sequences of arbitrary (comparable) values. knuth1997artv3
@@
We denote the bitwise-not operator by $\neg$. For example, $\neg 001_{2}=110_{2}$. Also, we can apply it to a set of outputs by flipping each output, e.g. $\neg\{000_{2},001_{2},010_{2}\}=\{111_{2},110_{2},101_{2}\}=\{101_{2},110_{2},111_{2}\}$. Also, we denote $\rho_{n}$ the reflection permutation on $[n]$, i.e. $\rho_{n}(i)=n-1-i$.
@@
bundala2014optimal uses this property to reduce the search space of the prefixes. We are giving a short proof here for completeness.
@@
In practice, it is not trivial to check if a network output set $A$ is isomorphic to the subset of another one $B$. We use two heuristics to speed up the check, one for positive cases, and one for negative cases.
Consider $A$ as an $n\times m_{A}$ Boolean matrix, where $n$ is the number of channels, $m_{A}$ is the number of outputs in $A$, and $A_{i,j}\in\{0,1\}$ is the output on $i$-th channel of the $j$-th output. Similarly, we consider $B$ as an $n\times m_{B}$ Boolean matrix. We sort the $n$ rows of $A$ and $B$ by the row sum, respectively. If multiple rows have the same sum, we permute them randomly. If the columns of $B$ covers the columns of $A$, then $A$ is isomorphic to a subset of $B$. We repeat this random process a couple of times. In practice, it detects more than $90\%$ of the redundant outputs efficiently.
To quickly filter negative cases, we compute the row sums and column sums of $A$ and $B$. Denote the row sum sequence of $A$ by $\mathrm{RowSumA}_{i}$ for $i=0,1,\ldots,n-1$, and the column sum sequence of $A$ by $\mathrm{ColSumA}_{j}$ for $j=0,1,\ldots,m_{A}-1$. Similarly, we have $\mathrm{RowSumB}$ and $\mathrm{ColSumB}$. We sort $\mathrm{RowSumA}$, $\mathrm{RowSumB}$, $\mathrm{ColSumA}$, and $\mathrm{ColSumB}$, respectively. If $A$ is isomorphic to a subset of $B$, then $\mathrm{RowSumA}_{i}\leq\mathrm{RowSumB}_{i}$ for all $i=0,1,\ldots,n-1$, and $\mathrm{ColSumA}$ is a subsequence of $\mathrm{ColSumB}$. This technique has been commonly used in isomorphism detections.
@@
3 Construction of a Sorting Network for 28 Channels with 13 Layers
@@ 3.1 High-Quality 16-Channel and 12-Channel 5-Layer Prefixes
We enumerate all non-redundant reflection-symmetric 12-channel 5-layer prefixes by extending and pruning one layer at a time.
@@ 3.4 Implementation and Running Time
We ran our program on a Mac mini M2 with 16GB RAM. The entire computation takes less than 20 minutes to find the solution shown in Figure 2. Specifically, finding the best 4 prefixes for 12 channels with 5 layers takes 12 minutes, while finding the best 2 prefixes for 16 channels with 5 layers takes 1 minute. We use MiniSat to solve the 8 SAT instances in parallel. It takes half a minute to find one solution, and 5 minutes to solve all 8 problems. All other steps complete very quickly.

### v2

@@ 1 Introduction
Our construction of n=28 and d=13 is based on ehlers2017merging. As we have more channels and one more layer, we use two new techniques to reduce the search space: assume the network is reflection-symmetric, and greedily extend the 24-channel prefix by one layer before solving the SAT problem. The full construction is found within 20 minutes on a common desktop computer.
@@ 2 Preliminaries and Redundant Prefixes
If a comparator network on $n$ channels sorts all $2^{n}$ Boolean sequences into non-decreasing order, then it correctly sorts all sequences of arbitrary (comparable) values. knuth1997artv3
@@
We denote the bitwise-not operator by $\neg$. For example, $\neg 001_{2}=110_{2}$. When it is applied to a set of outputs, it flips each output, e.g. $\neg\{000_{2},001_{2},010_{2}\}=\{111_{2},110_{2},101_{2}\}=\{101_{2},110_{2},111_{2}\}$. Also, For an integer $n$, we denote $[n]$ the set $\{0,1,\ldots,n-1\}$, and $\rho_{n}$ the reflection permutation on $[n]$, i.e. $\rho_{n}(i)=n-1-i$.
@@
Bundala and Závodnỳ bundala2014optimal use this property to reduce the search space of the prefixes. We are giving a short proof here for completeness.
@@
In practice, it is not trivial to check if a network output set $A$ is isomorphic to the subset of another one $B$. We use two techniques to speed up the check, one for positive cases, and one for negative cases.
Consider $A$ as an $m_{A}\times n$ Boolean matrix, and $A_{i,j}\in\{0,1\}$ is the output on $j$-th channel of the $i$-th output. Similarly, we consider $B$ as an $m_{B}\times n$ Boolean matrix. We sort the $n$ columns of $A$ and $B$ by the column sum, respectively. If multiple columns have the same sum, we permute them randomly. If the rows of $B$ covers the rows of $A$, then $A$ is isomorphic to a subset of $B$. We repeat this random process a couple of times. In practice, it detects more than $90\%$ of the redundant outputs efficiently.
To quickly filter negative cases, we compute the column sums and row sums of matrix $A$ and $B$. Denote the row sum sequence of $A$ by $\mathrm{RowSumA}_{i}$ for $i\in[m_{A}]$, and the column sum sequence of $A$ by $\mathrm{ColSumA}_{j}$ for $j\in[n]$. Similarly, we compute $\mathrm{RowSumB}$ and $\mathrm{ColSumB}$. We sort $\mathrm{RowSumA}$, $\mathrm{RowSumB}$, $\mathrm{ColSumA}$, and $\mathrm{ColSumB}$, respectively. If $A$ is isomorphic to a subset of $B$, then $\mathrm{ColSumA}_{j}\leq\mathrm{ColSumB}_{j}$ for all $j\in[n]$, and $\mathrm{RowSumA}$ is a subsequence of $\mathrm{RowSumB}$. This property is commonly used in bipartite graph isomorphism and embedding detections and sorting networks codish2016sorting.
@@
3 Construction of Depth-13 Sorting Network for 28 Channels
@@ 3.1 High-Quality 16-Channel and 12-Channel 5-Layer Prefixes
We enumerate all the non-redundant reflection-symmetric 12-channel 5-layer prefixes by the generate-and-prune approach codish2016sorting one layer at a time.
@@ 3.4 Implementation and Running Time
We ran our program on a Mac mini M2 with 16GB RAM. The entire computation takes less than 20 minutes to find the solution shown in Figure 2. Specifically, finding the best 4 prefixes for 12 channels with 5 layers takes 7 minutes, while finding the best 2 prefixes for 16 channels with 5 layers takes 1 minute. We use MiniSat to solve the 8 SAT instances in parallel. It takes half a minute to find one solution, and 5 minutes to solve all 8 problems. All other steps complete very quickly.